OpenStudy (anonymous):

Solve by taking the square roots of both sides. (x + 4)2 - 5 = 2 a.3,x=-11 b.4+7 squared ,x=-4+7 squared c. x=-4+7 squared , x=-4-7 squared d. x=-4+11 squared , x=-4-11 squared

OpenStudy (anonymous):

c. x=-4+7 squared , x=-4-7 squared

OpenStudy (anonymous):

thanks :)